OpenStudy (anonymous):

HELP MEEEEEEE PLEASE IM GIVING MEDALS Part 1: Create your own quadratic equation that cannot be solved by factoring, but can be solved using the quadratic formula. Identify the values of a, b, and c, and find the solutions using the quadratic formula. Show all work to receive credit. Part 2: Using complete sentences, explain how you know that the equation from Part 1 cannot be solved by factoring, but can be solved by using the quadratic formula.

OpenStudy (anonymous):

@amistre64 @shrutipande9 @JetWave600 @help123please. @Hero some one please help

OpenStudy (anonymous):

@AravindG

OpenStudy (anonymous):

In order to form an equation that doesn't factor in real numbers, choose a, b, and c so that b^2-4ac is negative. For example: b = 1, a = 1, c = 1: x^2 + x + 1 = 0 Using the quadratic formula: x = (-1 ± √(1 - 4)) / 2 x = (-1 ± √-3) / 2 x = (-1 ± i√3) / 2 where i^2 = -1 It cannot be factored in real numbers because the discriminant is less than zero. Straight off yahoo.

OpenStudy (anonymous):

i cant use that b/c it came off yahoo

OpenStudy (anonymous):

@genius12

OpenStudy (anonymous):

@Preetha